OpenStudy (anonymous):

n=4; 3 , 1/3 and 1+2i are zeros

OpenStudy (anonymous):

...

OpenStudy (dumbcow):

write it in factored form: however with complex roots, do the following x = 1 +-2i x-1 = +-2i (x-1)^2 = 4i^2 = -4 (x-1)^2 +4 = 0 x^2 -2x +5 = 0 --> (x-3)(x-1/3)(x^2 -2x +5) Now just multiply it out

OpenStudy (dumbcow):

also if you don't want to deal with fractions: x = 1/3 3x = 1 3x-1 = 0 Replace (x-1/3) with (3x-1) since they both yield the same zero of x=1/3

OpenStudy (anonymous):

ok thank you